OpenStudy (superhelp101):

what is the smallest value of (look below) for the interval 0 is less than or equal to theta is less than 2pie

OpenStudy (superhelp101):

\[2\sin^2(\theta)-\cos \theta-1=0\]

OpenStudy (superhelp101):

@Hero can u please help me

OpenStudy (irishboy123):

change the sin to cos using the simplest trig identity you can think of, and you have a quadratic in cos.